Page tree
Skip to end of metadata
Go to start of metadata
LTP4 Block STP cycleIOR delivery opensIOR deadlineSTP POR deadline (on SOC) Planned uplink by MOCMIBsStart of STP cycle onboardCommentsSOC ESA responsibleMain operator






2

158Wed 7 JulyWed 14 July (1200  )Tue 20 July (12:00)

Assumed Saturday 

on MIB from stp157

(22:00)

TCM type 2

IS, SPICE CMS and STIX ANALYSIS. SOLOHI will be off (not sending IORs) until STP160.

SCO-2
159Wed 7 JulyWed 21 July (1200  )Tue 27 July (12:00)

Assumed Saturday 


(19:00)

TCM type 2; SA ROTATION

IS, SPICE CMS (x2) and STIX ANALYSIS. SOLOHI will be off (not sending IORs) until next STP period.

160Wed 7 JulyTue 27 July (12:00) 

Friday 30 July  (12:00)

PORS  by 2 Aug latest

Assumed Tuesday 


(21:00)

TCM

IS, SPICE CMS, SOLOHI EXT window for venus observations (must be off though for VGAM, despite ext window overlapped with VGAM MAINT window – switches ON during 2021-08-07T14:58:00 EMC_PREF_NOISY).

161Wed 7 JulyThurs 29 July (08:00)Tue 3 Aug (12:00)

Assumed Saturday  aim for 6/8


  (15:15)

During VGAM, the RS instruments have to be switched off, this includes METIS, PHI, SOLOHI, STIX, SPICE OU and EUI OU. SPICE CMS IS OK.

SoloHI switches OFF 2021-08-09T04:00:00 before the VGAM 2 window

Note that following the VGAM, SWA HIS&EAS and SPICE OU cannot be switched back on until at least 1h after closest approach

SPICE deco heater commanding needed in HSD POR because of VGAM.

Outside of MAINT, STIX ANALYSIS and CMS



3

162

Wed 28 July


config update (SPICE/PHI) needed after window opening

Tue 3 Aug (1700) 

Later for PHI and SPICE

Wed 11 Aug (11:00 UTC)


Assumed Wednesday  

MIB (SPICE/PHI/STIX/SWA/HI) update on OPSLAN

aim for  MIB deployment after uplink of products for STP16) on 6/8 before the CEB pass starting at 09:12 UTC.  TBC depends on VGAM.

(06:00)

TCM, Roll manoeuvre, 

IS, Pre EXT window, SPICE CMS , SPICE functional (SPICE deco heater commanding needed in HSD POR  because SPICE HSD opens AND closes in this STP).

RS EXT windows lots of RS ops, also HSD operations and METIS TAC.

CFG needed after window opening


SCO-2

163

Wed 28 July


EUI MIB not needed until 164

Wed 11 Aug (12:00)

Later for EPD (EUI should not have IORs)

Tue 17 Aug (12:00)

Assumed Friday 

MIB (EUI) Update

Available 13th Aug, SOC can wait until 164

 

TCM

IS observations
RS observations for Metis, PHI and STIX
(EUI will not command the synoptics at start STP)
also HSD operations

Note deco heater may have to be turned off again because of SPICE observing.

PHI + EUI FSW update through PDORs, coordinated by MOC (23 and 25 Aug passes)

SCO-2




4

164Thursday 12 AugWed 18 Aug (12:00)Tue 24 Aug(12:00)

Assumed Friday(12:00)  


 

TCM, MAINT SA_RELUBRICATION

IS, RS EXT windows lots of RS ops, also HSD operations and RPW TAC

Jayne Lefort SCO-2
165Thursday 12 AugWed 25 Aug (12:00)Tue 31 Aug(12:00)

Assumed  Friday 


 

MAINT ANY

IS, RS /RS EXT windows lots of RS ops, also HSD operations. PHI TCA, SPICE TAC, EUI TAC,

166Thursday 12 AugFri 3 Sep (12:00)Thurs 9 Sep(12:00)

Assumed Monday  


 

IS, RS windows lots of RS ops, also HSD operations. VTSP updates
167Thursday 12 AugWed 8 Sep (12:00) Tue 14 Sep(12:00)

Assumed Friday 


  (23:00)

IS, RS  EXT windows lots of RS ops, also HSD operations. everyone

All dates in 2021, UTC


MIB version to be used: 

Overview of all MIB versions and applicability: MIB and FOP


  • No labels